### v3

# EventHallusion: Diagnosing Event Hallucinations in Video LLMs

Abstract
Recently, Multimodal Large Language Models (MLLMs) have made significant progress in the video comprehension field. Despite remarkable content reasoning and instruction following capabilities they demonstrated, the hallucination problem of these VideoLLMs is less explored compared with its counterpart in the image domain. To mitigate this gap, we propose EventHallusion, a novel benchmark that focuses on assessing the VideoLLMs’ hallucination toward event, the crux of video analysis. From a hallucination attribution perspective, our EventHallusion benchmark is curated to assess a VideoLLM’s susceptibility toward language priors and vision-language biases. On the other hand, we also propose a simple yet effective method, called Temporal Contrastive Decoding (TCD), to tackle the hallucination problems of VideoLLMs. The proposed TCD method rectifies the model’s bias toward its priors during the decoding stage by comparing the original video with a modified version, in which temporal cues are disrupted. Through comprehensive evaluation of eight open-source and two closed-source VideoLLMs on the proposed EventHallusion benchmark, we observe that the open-source models suffer significantly from hallucination problems, whereas the closed-source ones perform markedly better. By further equipping open-source VideoLLMs with the proposed TCD approach, evident performance improvements are achieved across most metrics in the EventHallusion benchmark. Our codes and benchmark data are available at https://github.com/Stevetich/EventHallusion.
Figure 1: Illustration of two types of underlying reasons that may lead to existing VideoLLMs’ hallucinations. We use red to indicate the hallucinations in answers generated by existing state-of-the-art VideoLLMs, namely Video-LLaVA and ShareGPT4Video.
1 Introduction
The field of deep learning has witnessed the extraordinary advancement of Large Language Models (LLMs) [6, 36, 10] and Multimodal Large Language Models (MLLMs) [4, 28, 5, 15, 13]. Benefiting from dramatic semantic reasoning and instruction-following capabilities, LLMs and MLLMs have become versatile assistants across multiple disciplines  [13, 41]. Despite their great success, they are still trapped in the hallucination problem, where models generate plausible but factually incorrect responses [18], potentially threatening reliability and security in applying them in real-world scenarios [33, 14]. To systematically evaluate hallucinations of MLLMs, a thread of benchmarks has been developed.  [22, 11, 29]. Specifically, POPE examines dataset distributions to investigate the hallucination phenomenon in MLLMs, focusing on frequently occurring objects and common object combinations, while MME and MMBench assess hallucinations through systematic evaluations of object existence, attributes, and relations.
However, these benchmarks primarily assess hallucinations within static scenes, such as those involving objects, attributes, and inter-object relations, and therefore do not effectively address the needs of emerging VideoLLMs  [31, 23]. Two recent benchmarks [37, 12] further extend their evaluation to dynamic scenes with event-related queries. In pursuit of the rigorous evaluation, both of them curate more challenging event comprehension tasks. Specifically, HallusionBench [12] reverses all frames within a video, subsequently querying the model to determine whether the original event occurs in the altered video. VideoHallucer [37] selects long videos containing multiple events and then queries the model to evaluate the chronological order of these events. Although advancing the evaluation of event hallucination, they still suffer from the following limitations. On the one hand, the frame-reversed video in HallusionBench can deviate from the distribution of natural videos, as not all events remain meaningful when played in reverse. On the other hand, merely judging the chronological order of events given in the questions fails to capture the model’s authentic ability to comprehend the detailed content of the events [^1] A more detailed analysis of event-related VQA samples in HallusionBench [12] and VideoHallucer [37] is provided in the supplementary materials due to page limitations..
When delving into VideoLLMs’ event-related hallucinations, we identify two primary causes rooted in the development processes of prevalent VideoLLMs. Firstly, building upon the foundation of LLMs, VideoLLMs inherently inherit their language priors [3, 2], which manifests as an overconfidence in their language knowledge. As illustrated in the left column of Fig.1, both Video-LLaVA and ShareGPT4Video are easily misled by the frequently occurring but misleading event “person drink coffee” asked in the question. Moreover, due to the high cost associated with scaling up high-quality video-language data, prevalent VideoLLMs typically acquire multimodal knowledge from large-scale image-language datasets. However, in these datasets, specious event-related captions often arise based solely on a single image, which inevitably introduces spurious scene-event correlations. As shown in the right column of Fig.1, both Video-LLaVA and ShareGPT4Video intuitively make the conclusion of “cats are playing” when perceiving two cats within the video, without respecting the truth.
Toward this end, in this paper, we propose EventHallusion, a novel benchmark that systematically evaluates event-related hallucinations of state-of-the-art VideoLLMs. From a hallucination attribution perspective, our EventHallusion is curated to assess the susceptibility of a VideoLLM to language priors and vision-language correlation biases, respectively. Specifically, our benchmark consists of two basic categories as shown in Fig.2: (1) Susceptibility to language priors: In this category, all questions are annotated with a misleading event designed to prompt the model to make decisions based on its language priors. Within such events, while objects align with those depicted in the video, their actions are deliberately misrepresented. (2) Susceptibility to vision-language correlation biases: In this category, all videos are meticulously selected to ensure that the included events are infrequent yet plausible. When presented with these infrequent events, the model can easily be misled to interpret the event as one that frequently co-occurs with the objects or scenes depicted in the video, based on its memory. It is worth mentioning that through comprehensive evaluation, we find that existing open-source VideoLLMs suffer from severe event hallucination issues on our benchmark, which could offer promising avenues for enhancing the VideoLLM’s capabilities in the future research.
Figure 2: Illustration of examples in our EventHallusion benchmark. Questions are basically divided into two categories: Susceptibility to Language Prior and Susceptibility to Vision-Language Correlation Bias. The contents that induces VideoLLMs to generate event-related hallucinations are marked by red.
Moreover, we propose a simple yet effective method called Temporal Contrastive Decoding (TCD) to reduce the event-related hallucinations of existing VideoLLMs. Without any additional training budgets, our TCD rectifies the model’s bias to its prior during the decoding stage. Technically, for each given video, we first apply the temporal downsampling operation to create a modified counterpart. Both the original video and its modified counterpart are then input into the VideoLLMs. During the decoding stage, the logits prediction of the original video is adjusted by the logits prediction of the counterpart. The rectified logits are then used to sample the next token and continue the decoding process. Within the proposed TCD, the modified counterpart disrupts the temporal cues while preserving spatial object and scene information, making its predicted logits more susceptible to bias from the model’s event-related prior knowledge. Subtracting these amplified priors in the logits space during the decoding process helps guide the model to make more reliable judgments. As a plug-and-play approach, our TCD can consistently boost existing VideoLLM’s performances on the proposed EvenHallusion as shown in Tab.2.
In summary, our contributions are threefold: (1) We propose EventHallusion, a novel video QA benchmark focused on assessing event-related hallucinations in VideoLLMs; (2) We introduce Temporal Contrastive Decoding (TCD), a simple yet effective approach that can reduce hallucinations across a wide range of VideoLLMs in a plug-and-play manner; (3) We also conduct comprehensive ablation studies to provide more in-depth analysis.
2 Related Works
Video Large Language Models. In the natural language processing field, Large Language Models (LLMs) have demonstrated astounding performance in solving traditional tasks. What’s more, LLMs also exhibit extraordinary generalization in handling self-defined tasks described by user’s instructions, making them capable of playing the role of versatile assistants. Benefiting from such abilities, Multimodal Language Models (MLLMs) have emerged by introducing visual inputs into LLMs, promoting the unified comprehension of both textual and visual content.
Video-supported MLLMs are further proposed since the established vision-text understanding ability of image MLLMs. To enable MLLMs to comprehend video content and the concept of temporality, numerous methods have been put forward. Video-ChatGPT [32] utilizes spatial and temporal pooling operations to conduct video modeling, while Video-LLaMA [43] introduces a Video Q-Former to aggregate frame-level representations. Vista-LLaMA [31] preserves the equal distance between all visual tokens and any language tokens and proposes the temporal Q-Former projector to further enhance the representations of the video and the modeling of temporality. LLaMA-VID [23] proposes a dual-token paradigm, using a context token and a content token to represent each frame, which effectively supports the comprehension of long videos and maintains temporal information. However, despite their excellent performance on existing benchmarks, these methods still suffer the hallucination problem in temporality.
Video Evaluation Benchmarks. Despite the extraordinary advancement of MLLMs, it is of great significance to conduct quantitative evaluations of existing models. To thoroughly evaluate the performance of image MLLMs, a wide scope of efforts has been made [11, 29, 24]. Their evaluation comprises various aspects, including perception, reasoning, and cognition. With the development of VideoLLMs, several benchmarks have also been proposed to assess video comprehension capacities. Tempcompass [30] focuses on the thorough evaluation of the temporal perception ability of VideoLLMs, including action, speed, direction, attribute change, and event order. FunQA [38] proposes utilizing humorous, creative, and magic videos to assess the in-depth comprehension ability of existing VideoLLMs.
Despite previous works sharing similarities with ours, there exist several critical differences. First, we focus on exploring and evaluating the hallucination phenomenon in the event aspects of existing VideoLLMs. Furthermore, we emphasize and validate that the event hallucination phenomenon of video LLMs is influenced by language prior. Second, based on this conclusion, we collect videos that are inclined to cause event hallucinations in VideoLLMs and construct the benchmark for evaluation. Third, we propose a Temporal Contrastive Decoding (TCD) to alleviate the event hallucination, which exhibits effectiveness in multiple VideoLLMs.
Hallucination in MLLMs. Hallucination is a phenomenon widely observed in MLLMs. The occurrence of hallucinations is attributed to various factors, including imbalanced training data distribution, weak attention to visual content, and other possible reasons. To alleviate the impact of hallucinations in image MLLMs, several methods have been proposed. Woodpecker [40] adopts a post-hoc correction mechanism that utilizes additional extracted visual cues to rectify previously generated responses, while Volcano [17] proposes a critique-revise-decide process to self-revise their initial generated answers. Besides, LRV-Instruction [27] and HalluciDoctor [42] propose to mitigate this problem from the data perspective to provide more balanced, high-quality instruction tuning datasets.
Some other works propose to alleviate hallucination during the inference phase, in which the contrastive decoding (CD) method has been widely adopted due to its excellent performance and the absence of additional training. VCD [18] employs the original and distorted images for comparison in contrastive decoding, thereby reducing the impact of statistical bias and language priors, while IBD [44] constructs the comparison pair by introducing the image-biased and original MLLMs. Despite the extensive exploration of adopting CD in ImageLLMs, there is still a lack of effort in employing CD to handle hallucinations in VideoLLMs. Inspired by this, we propose Temporal Contrastive Decoding (TCD) to mitigate event hallucinations in VideoLLMs.
3 EventHallusion Benchmark
In this section, we first elaborate on the data curation process of two basic categories of our EventHallusion benchmark. Afterward, we demonstrate the distribution statistics of the whole EventHallusion benchmark. Finally, we introduce the specific evaluation protocol of different question types.
3.1 Data Curation
Susceptibility to language priors. In this category, we first collect videos describing the events that occur in daily life from the internet. To avoid information leakage, we blur the text captions printed on videos, thereby encouraging models to analyze events solely based on visual content. However, when annotating the questions, we explicitly include another frequently occurring event that is inconsistent with the video content to trigger the language priors within the LLM. With these “misleading triggers”, we formulate questions as a binary choice problem as shown below:
“Did the [subject] [event] happen in the video?”
Here, the combination of [subject] and [event] correspond to the “misleading triggers” aforementioned. In Fig.2, we refer to this type of sample as “Common Events with Misleading Questions”.
Susceptibility to vision-language correlation biases. In this category, the primary challenge lies in sourcing the appropriate video data. Since existing publicly available video datasets (e.g, ActivityNet [7], [9], [34]) collect videos from daily life, the events they include cater to the vision-language correlation biases stored in VideoLLMs, resulting in high recognition precision. Therefore, instead of refactoring existing video datasets, we meticulously collect videos from prevalent Internet video platforms and filter them with rigorous cross-validation. Specifically, we begin by manually collecting videos that depict infrequent events based on human consensus. To minimize personal bias, we engage three annotators to assess whether the events in these selected videos are indeed uncommon. Only videos that receive unanimous positive judgments from all annotators are retained as testing samples. Similar to the first category, we also employ the blurring operation on videos to prevent information leakage. We further divide these videos into two groups: (1) Videos where infrequent events occur throughout the entire duration, referred to as “Entire Rare Events” in Fig.2. (2) Videos containing a mix of both frequent and infrequent events, referred to as “Mix Common-Rare Events” in Fig.2.
For constructing questions, we design both binary and open-ended types for comprehensive evaluation. Specifically, the binary questions ask the model whether a plausible but incorrect common event occurs in the given video. The open-ended questions, on the other hand, simply prompt the model to freely describe the video content. Formally, binary and open-ended questions for the “Entire” category can be formulated as presented below:
Binary: “Did the [subject] [event] in the video?”,
Open-Ended: “Please describe this video in detail.”,
For the “Mix” category, since both common and rare events can happen in the video, we reformulate the binary questions as below:
“Did the [subject] [event] throughout the entire video?”,
“Did any accident or anything unexpected happen in the video?”,
Dataset distribution statistics. In total, our EventHallusion benchmark comprises 400 videos and 711 questions. To ensure comprehensive coverage across diverse domains, these videos span seven scenarios, including Pet&Animal, Sports Competition, Food&Drink, Gym Exercises, Vehicle, Life Record, Nature, as shown in Fig.3. The distribution of video lengths in our benchmark is presented in Fig.4, showing that most videos are relatively short. This is because longer videos often involve significant scene fluctuations, which introduce additional challenges for VideoLLMs beyond the uni-modal and multi-modal biases we aim to examine in this paper.
Figure 3: Video category and questions distribution. “Entire” implies that the entire video depicts the rare event. “Mix” implies that common and rare events interleavedly occur in the video. “Misleading” implies the video with the common event, but with its question customized to include misleading event demonstration.
Figure 4: The distribution of video length within our EventHallusion benchmark.
3.2 Evaluation Protocol
Binary question answering. Following previous benchmarks [37, 22], we compare the first word of the model’s answer with the ground-truth answer to determine whether the model answered the question correctly. This approach is justified by the VideoLLM’s exceptional instruction-following capability, which typically enables it to answer binary choice questions with “yes” or “no” at the beginning. We calculate the accuracy as the evaluation metric for binary questions.
Open-ended question answering. For automatic evaluation, we feed both generated and ground-truth answers into GPT-4o, using it as a knowledgeable evaluator to determine whether the model answers the question correctly following prior benchmarks [30]. We also use accuracy as the evaluation metric. Detailed templates of prompting GPT-4o are provided in the supplementary materials due to page limitation.
Figure 5: Framework of Temporal Contrastive Decoding. We construct the temporal modeling distorted video as the comparison group with fewer frames extracted and then contrast their logits to mitigate the event hallucination.
4 Temporal Contrastive Decoding
To reduce the aforementioned hallucinations in VideoLLMs, the most straightforward approach is to further enrich debiased data for finetuning. However, this approach is particularly resource-intensive for collecting such videos and annotating QA pairs. To address this, we present our Temporal Contrastive Decoding (TCD) method, which effectively alleviates hallucinations in videos without requiring additional training efforts.
Motivated by the previous work [18], our TCD method aims to reduce event hallucinations by adjusting the distribution of predicted logits at each decoding step, using a contrastive counterpart that is more aligned with the model’s inherent biases. For clarity, we first review the standard auto-regressive decoding process of a VideoLLM. Given the text instruction $x$ and the video $V$, the predicted logit $z_{t}^{ori}$ at the $t$-th timestep can be formulated as:
$z_{t}^{ori}=\operatorname{logit}_{\theta}\left(y_{t}\mid V,x,y_{<t}\right)$ (1)
where $y_{t}$ is the token to be generated at time step $t$, and $y_{<t}$ denotes the generated token sequence before time step $t$. $y_{t}$ is based on $x$, $V$, and all previously generated tokens $y_{<t}$ by selecting the token with the highest probability.
When constructing a contrastive counterpart, we chronologically downsample the original video as shown in the Fig.5, aiming to weaken the temporal cues while emphasizing object cues. Therefore, the contrastive counterpart’s probability prediction $z_{t}^{con}$ at timestep $t$ can be formulated as:
$z_{t}^{con}=\operatorname{logit}_{\theta}\left(y_{t}\mid S(V),x,y_{<t}\right)$ (2)
where $S(\cdot)$ represents the temporal downsampling operation. With the generated $z_{t}^{ori}$ and $z_{t}^{con}$, the modulated probability distribution $p_{t}$ at the current timestep can be calculated as:
$p_{t}=\operatorname{softmax}\Big{[}(1+\alpha)z_{t}^{ori}-\alpha z_{t}^{con}\Big{]}$ (3)
where $\alpha$ is the hyperparamter to control the contrastive strength during the decoding. However, applying such a contrastive process throughout the entire decoding stage would lead to significant model degradation, as it is not always consistent for $z_{t}^{ori}$ and $z_{t}^{con}$ to have mismatched distribution. When $z_{t}^{ori}$ and $z_{t}^{con}$ are consistent and correct, employing Eq.3 to modulate the probability impairs the model’s judgment. Therefore, we introduce another hyperparameter $\beta$ to constrain the token set used for TCD. The motivation behind this design is twofold: when the model exhibits high confidence in a specific token, we conduct TCD on a smaller subset of tokens with similarly high confidence. Conversely, when the model’s confidence is distributed more evenly across tokens, we apply a lower threshold, enabling TCD on a broader range of tokens.
Table 1: Performance comparison of VideoLLMs on EnventHallusion. We exhibit the accuracy of both open-sourced and closed-sourced models on binary classification task and detailed description matching task on our three types of videos respectively. We use red to highlight the best results in close-source VideoLLMs and blue for open-source VideoLLMs. In addition, we show the overall success rate of each model on our binary classification task that answers ”yes” or ”no” in the first word.
Models Entire Mix Misleading Overall
Binary Desc. Binary Desc. Binary Binary Desc.
Open-source Models
VideoChatGPT [32] 14.91 5.50 56.99 3.63 21.57 36.43 4.30
VideoChat2 [21] 16.67 4.59 47.67 1.55 22.55 32.76 2.64
PLLaVA [39] 45.61 16.51 58.55 3.11 81.37 60.64 6.05
LLaMA-VID [23] 30.70 16.51 73.58 7.77 43.14 54.03 10.90
LLaVA-NeXT-Video [19] 49.12 9.17 70.47 4.66 71.57 64.79 6.29
VILA [26] 53.51 20.18 58.55 17.62 83.33 63.33 18.54
ShareGPT4Video [8] 11.40 0.00 67.88 5.18 6.86 49.14 9.82
Video-LLaVA [25] 30.70 8.26 57.51 7.25 41.18 45.97 7.62
Close-source Models
GPT-4V [1] 57.89 11.01 79.27 16.58 92.16 76.53 14.57
Gemini-1.5-Pro [35] 61.40 49.54 83.42 39.90 96.08 80.44 43.38
GPT-4o [1] 83.33 58.72 92.75 41.97 100.00 91.93 48.01
Table 2: Results of TCD on EventHallusion. Default denotes the standard decoding strategy, while TCD represents our Temporal Contrastive Decoding method. We exhibit the accuracy of the binary classification task. The best results of each category are bolded.
Model Decoding Entire Mix Misleading Overall
LLaVA-NeXT-Video [19] Default 49.12 70.47 71.57 64.79
TCD 50.88 73.58 74.51 67.48
VideoChat2 [21] Default 16.67 47.67 22.55 32.76
TCD 17.54 49.74 25.49 34.72
VILA [26] Default 53.51 58.55 83.33 63.33
TCD 50.88 66.32 84.31 66.50
5 Experiments
In this section, we first evaluate multiple prevalent VideoLLMs on our EventHallusion, including both open-sourced and private MLLMs. Subsequently, we examine the effectiveness of our proposed TCD, and also discuss its more specific design choices.
5.1 Evaluation on EventHallusion
5.1.1 Selected VideoLLMs for Evaluation
We test 8 open-sourced Video and 2 close-sourced VideoLLMs on our EventHallusion. Open-sourced VideoLLMs comprise VideoChatGPT [32], VideoChat2 [21], PLLaVA [39], LLaMA-VID [23], VideoLaVIT [16], LLaVA-NeXT-Video [19], VILA [26], ShareGPT4Video [8], and Video-LLaVA [25]. For close-sourced VideoLLMs, we select the three most popular video-supported MLLMs, namely GPT-4V [1], GPT-4o and Gemini-1.5-Pro [35], to conduct the systematic comparison of the robustness between open-sourced and close-sourced MLLMs.
For the inference setting of each VideoLLM, we follow their original setting including conversation mode and other hyperparameters. For the number of frames, we sample at least 12 frames for each VideoLLM to ensure the complete preservation of event information. In addition, to ensure the stability and reproduction of our results, we adopt the greedy search as the default decoding strategy for all open-sourced models. For the close-sourced models, we sample 8 frames.
5.1.2 Performance Analysis
We exhibit the performance of each VideoLLM in Tab 1 on the binary classification and detailed description matching task. As the results indicate, open-source VideoLLMs exhibit a consistent weakness in event hallucinations. In contrast, closed-source VideoLLMs demonstrate greater robustness against event hallucinations across all categories. A deeper analysis is provided below: (1) In evaluating the susceptibility of VideoLLMs to vision-language correlation biases (i.e, “Entire” and “Mix” in Tab.1), most VideoLLMs perform worse when confronted with videos containing a mix of common and rare events, compared to those containing only rare events. This result indicates that frequent events can distract the model’s attention, imposing higher challenges to the model. (2) In evaluating the susceptibility of VideoLLMs to language priors (i.e., “Misleading” in Tab.1), closed-source models demonstrate overwhelming advantages compared to open-source models, suggesting that these closed-source VideoLLMs are built upon language foundation models that are significantly more powerful than their open-source counterparts.
5.2 TCD Evaluation
5.2.1 VideoLLM Baselines.
To evaluate the effectiveness of our TCD in addressing event hallucinations, we adopt LLaVA-NeXT-Video [19], VideoChat2 [21], and VILA [26] as the foundation backbones. LLaVA-NeXT-Video extends the ability of LLaVA [28] based on the projection layer design, while VideoChat2 adopts the Q-Former [20] to bridge the gap between videos and LLM. VILA, on the other hand, employs enhanced visual-language pre-training for better video comprehension. Based on these MLLMs, we explore the performance of TCD across different VideoLLM architectures and designs on EventHallusion.
5.2.2 Implementation Details.
We adopt different numbers of extracted frames to construct the original and distorted visual input: 32 and 8 frames for LLaVA-NeXT-Video, 12 and 8 frames for VILA, and 16 and 4 frames for VideoChat2. We use greedy search as the basic decoding paradigm.
5.2.3 Main Results Analysis
We exhibit the results of the binary classification task on EventHallusion in Tab.2. As shown in Tab.2, our TCD demonstrates significant performance across all three VideoLLMs, with increases in overall accuracy, confirming the effectiveness of TCD in mitigating event hallucinations on VideoLLMs with different architectures and visual-language pertaining paradigms. However, while TCD shows consistent and notable improvements on Mix and Misleading across all three models, the improvement on Entire is minimal, with even a decrease situation observed. This result suggests that entire rare events are a more challenging and hallucination-prone scenario for VideoLLMs.
5.2.4 Ablations on Contrastive Sample Constructions.
We also discuss some different alternative strategies when constructing the counterpart of the original video for performing our TCD method as shown in Tab.3. Specifically, we compare our base TCD method with adding Gaussian noises to video frames when employing TCD method (denoted as TCD+noises). As results have shown in Tab.3, While both TCD and TCD+noises demonstrate improvement on regular decoding. However, TCD+noises does not exhibit large improvements on simply using TCD method.
Table 3: Ablation study of different strategies when constructing the contrastive counterpart of the original video. The best results of each category are bolded.
Method Entire Mix Misleading Overall
Baseline 49.12 70.47 71.57 64.79
TCD 50.88 73.58 74.50 67.48
TCD+noise 51.75 73.58 74.50 67.73
Table 4: Ablation study of $\alpha$ in our TCD. The best results of each category are bolded.
$\alpha$ Entire Mix Misleading Overall
0.25 50.00 73.56 73.53 66.99
0.5 50.88 73.56 74.51 67.48
0.75 50.88 71.50 74.51 66.50
1.0 49.12 64.25 74.51 62.59
Table 5: Ablation study of $\beta$ in our TCD. The best results of each category are bolded.
$\beta$ Entire Mix Misleading Overall
0.1 50.88 73.58 74.51 67.48
0.2 50.88 73.58 74.51 67.48
0.5 50.88 74.09 74.51 67.73
0.5 50.00 74.61 74.51 67.73
5.2.5 Ablations on Hyper-parameters of TCD
We conduct ablation studies to investigate the effects of hyper-parameters $\alpha$ and $\beta$ on our TCD method as shown in Tab.4 and Tab.5 For the first hyper-parameter $\alpha$, with the growth of $\alpha$, TCD applies a more fierce contrast between normal and distorted video input. As results are shown in Tab.4, when $\alpha$ is 0.5, TCD exhibits best performance. While for $\beta$, the threshold used to control the token set for TCD does not exhibit a strong impact on the performance of TCD on our EventHallusion Benchmark.
5.2.6 Effect of Sampled Frames in Distorted Video.
We further explore the effect of sampled frames in distorted video on our TCD. Specifically, we adopt LLaVA-NeXT-Video as the backbone model with 32 frames sampled for the original video input and adjust the numbers of sampled frames in the distorted video. We conduct this experiment on the binary classification task of EventHallusion. As results have shown in Tab.6, TCD maintains a steady improvement in the range of 4 to 16 sampled frames. However, when the number of sampled frames in distorted video decreases from 4 to 2, the performance of TCD deteriorates. We attribute this phenomenon to the excessive destruction of the temporal modeling in the VideoLLM with too few frames incorporated. This leads to drastic changes in the logit of distorted video, thereby affecting the contrastive decoding performance.
Table 6: Ablation study of extracted video frames in the distorted group. The best results of each category are bolded.
Frames Entire Mix Misleading Overall
1 48.26 68.39 74.51 64.30
4 50.00 73.06 76.47 67.48
8 50.88 73.58 74.51 67.48
16 50.88 73.06 74.51 66.75
Figure 6: Qualitative Results of the TCD method on our EventHallusion Benchmark.
5.3 Qualitative Results
We exhibit an example of our question in EventHallusion. As shown in Fig.6, VideoLLM can be affected and generate hallucination answers, while our TCD method can alleviate this problem. Due to the page limit, we provide more examples in our suuplementary materials.
6 Conclusion
In this paper, we systematically investigated the event-related hallucination problem of existing VideoLLMs. We curated EventHallusion, which consists of multi-domain videos and annotated questions, both of which are specifically designed to trigger the priors encapsulated in the VideoLLMs themselves. In addition, we also proposed a simple yet effective method, called Temporal Contrastive Decoding (TCD), to alleviate the event-related hallucination problems.
Supplementary Material
1 Comparison with other video hallcination benmarks.
We exhibit the comparison between HallusionBench, VideoHallucer, and our EventHallusion in Figure 7. As shown in Figure 7, compared to HallusionBench and Videohallucer emphasize assessing the models’ sensitivity to events’ chronological orders, our proposed EventHallusion focuses on analyzing the event-related language prior and vision-language correlation bias encapsulated in existing VideoLLMs.
2 Automatic Evaluation For Detailed Description Matching
We apply GPT-4o to evaluate the detailed description matching task of our Entire, Mix, and Misleading categories respectively. For Entire, we prompt GPT-4o to judge if VideoLLMs’ detailed description of the video is consistent with the ground-truth rare event. And for Mix, we prompt GPT4o to focus on the occurrence of the rare-common mixed event. Finally, for Misleading, we prompt GPT4o to assess the consistency of the ground-truth event and search for any unrelated events. In addition, we set up the default system prompt to better guide GPT-4o’s evaluation, which is shown in Figure 8. The detailed prompts are shown in Figure  9, 10, 11, respectively.
Figure 7: Comparison between HallusionBench, VideoHallucer, and our EventHallusion. We use Green to highlight the content of the original video/question, and red to denote refactored video/question content to assess the model’s event-related hallucinations.
Figure 8: System prompt used in all automatic evaluation.
3 Qualitative Examples
3.1 Data examples of EventHallusion.
We present examples of Entire, Mix, and Misleading in Figure 12, 13, 14, incorporating responses of VideoLLMs. As shown in these examples, VideoLLMs are prone to generate hallucinated answers under all of our settings, indicating their deficiency in addressing event hallucinations.
3.2 Examples of GPT-4o’s Evaluation.
We present examples of GPT-4o’s evaluation of detailed description matching task on Entire, Mix, and Misleading respectively in Figure 15, 16, 17. As shown in these examples, GPT-4o have the capacity to judge the accuracy of video events.
3.3 Examples of TCD.
We present examples of our TCD in mitigating video event hallucinations. Specifically, we exhibit the generated response with and without TCD for LLaVA-NeXT-Video and VideoChat2 on our EventHallusion in Figure 18 and 19.
Figure 9: Prompt used for detailed description matching in Entire.
Figure 10: Prompt used for detailed description matching in Mix.
Figure 11: Prompt used for detailed description matching in Misleading.
Figure 12: Examples of Entire with responses of VideoLLMs.
Figure 13: Examples of Mix with responses of VideoLLMs.
Figure 14: Examples of Misleading with responses of VideoLLMs.
Figure 15: Examples of the evaluation of GPT-4o on the detailed description of VideoLLMs of Entire.
Figure 16: Examples of the evaluation of GPT-4o on the detailed description of VideoLLMs of Mix.
Figure 17: Examples of the evaluation of GPT-4o on the detailed description of VideoLLMs of Misleading.
Figure 18: Examples of TCD with LLaVA-NeXT-Video on our EventHallusion.
Figure 19: Examples of TCD with VideoChat2 on our EventHallusion.