### v6

# EventHallusion: Diagnosing Event Hallucinations in Video LLMs

Abstract
Recently, Multimodal Large Language Models (MLLMs) have made significant progress in the video comprehension field. Despite remarkable content reasoning and instruction following capabilities they demonstrated, the hallucination problem of these VideoLLMs is less explored compared with its counterpart in the image domain. To mitigate this gap, we propose EventHallusion, a novel benchmark that focuses on assessing the VideoLLMs’ hallucination toward event, the crux of video analysis. From a hallucination attribution perspective, our EventHallusion benchmark is curated to assess a VideoLLM’s susceptibility toward language priors and vision-language biases. On the other hand, we also propose a simple yet effective method, called Temporal Contrastive Decoding (TCD), to tackle the hallucination problems of VideoLLMs. The proposed TCD method rectifies the model’s bias toward its priors during the decoding stage by comparing the original video with a modified version, in which temporal cues are made ambiguous. Through the comprehensive evaluation of eight open-source and three closed-source VideoLLMs on the proposed EventHallusion benchmark, we observe that the open-source models suffer significantly from hallucination problems, whereas the closed-source ones perform markedly better. By further equipping open-source VideoLLMs with the proposed TCD approach, evident performance improvements are achieved across most metrics in the EventHallusion benchmark. Our codes and benchmark data are available at https://github.com/Stevetich/EventHallusion.
Figure 1: Illustration of two types of underlying reasons that may lead to existing VideoLLMs’ hallucinations. We use red to indicate the hallucinations in answers generated by existing state-of-the-art VideoLLMs, namely Video-LLaVA and ShareGPT4Video.
Introduction
The field of deep learning has witnessed the extraordinary advancement of Large Language Models (LLMs) (brown2020language; touvron2023llama; dettmers2024qlora) and Multimodal Large Language Models (MLLMs) (alayrac2022flamingo; liu2024visual; bai2023qwen; jiao2024lumen; jiao2024rode). Benefiting from dramatic semantic reasoning and instruction-following capabilities, LLMs and MLLMs have become versatile assistants across multiple disciplines  (jiao2024rode; yin2023foodlmm). Despite their great success, they are still trapped in the hallucination problem, where models generate plausible but factually incorrect responses (leng2024vcd), potentially threatening reliability and security in applying them in real-world scenarios (qian2024nuscenes; jiao2022more). To systematically evaluate hallucinations of ImageLLMs, a thread of benchmarks has been developed (pope; fu2023mme; liu2023mmbench). Specifically, POPE (pope) investigates the hallucination by mining the frequently occurring objects and their combinations. MME (fu2023mme) and MMBench (liu2023mmbench) evaluate hallucinations of the existence, attributes, and relations of objects.
In fact, based on the foundation of ImageLLMs, VideoLLMs have shown significant proficiency in understanding the content of single-frame images (wang2024videohallucer). However, their effectiveness diminishes when tasked with comprehending the event, namely the actions performed by humans or objects spanning the entire video (du2024towards; zelnik2001event; krishna2017dense). This limitation is further proved by our results presented in Fig.6. To further evaluate the event-oriented hallucinations of VideoLLMs, two recent benchmarks (wang2024videohallucer; guan2024hallusionbench) develop distinct event comprehension tasks. Specifically, HallusionBench (guan2024hallusionbench) reverses all frames within a video, subsequently querying the model to determine whether the original event occurs in the altered video. VideoHallucer (wang2024videohallucer) selects long videos containing multiple events and then queries the model to evaluate their chronological orders. Although advancing the evaluation of event hallucination, they still suffer from the following limitations. On the one hand, the frame-reversed video in HallusionBench can deviate from the distribution of natural videos, as not all events remain meaningful when played in reverse. On the other hand, merely judging the chronological order of events given in the questions fails to probe the model’s authentic ability to comprehend the complex events [^1] A more detailed analysis of event-related VQA samples in HallusionBench (guan2024hallusionbench) and VideoHallucer (wang2024videohallucer) is provided in the supplementary materials due to page limitations..
When delving into event-related hallucinations in VideoLLMs, we observe that such errors originate from a consistent issue across different training stages: imbalanced training data distributions, which inject prior knowledge and bias into the model. Initially, building upon LLM backbones, VideoLLMs memorize events with higher frequencies purely through textual statistics, thereby forming strong language-driven priors (agrawal2016analyzing; agarwal2020towards). For instance, the commonly described event “person drinks coffee” frequently appears in textual corpora, establishing a misleading prior that persists regardless of visual evidence (Fig.1, left). Moreover, due to the high cost associated with scaling up high-quality video-language data, prevalent VideoLLMs typically acquire multimodal knowledge from large-scale image-language datasets. However, similar imbalances surface during the vision-language alignment stage, where frequently co-occurring visual-textual pairs reinforce spurious object-event associations. As shown in the right column of Fig.1, both Video-LLaVA (videollava) and ShareGPT4Video (chen2024sharegpt4video) intuitively conclude “people are riding the bicycle” simply from the presence of a person and a bicycle in the scene, disregarding the actual video context.
Toward this end, in this paper, we propose EventHallusion, a novel benchmark that systematically evaluates event-related hallucinations of state-of-the-art VideoLLMs. From a hallucination attribution perspective, our EventHallusion is curated to assess the susceptibility of a VideoLLM to language priors and vision-language correlation biases, respectively. Specifically, our benchmark consists of two basic categories as shown in Fig.2: (1) Susceptibility to language priors: In this category, all questions are annotated with a misleading event designed to prompt the model to make decisions based on its language priors. Within such events, while objects align with those depicted in the video, their actions are deliberately misrepresented. (2) Susceptibility to vision-language correlation biases: In this category, all videos are meticulously selected to ensure that the included events are infrequent yet plausible. When presented with these infrequent events, the model can easily be misled to interpret the event as one that frequently co-occurs with the objects or scenes depicted in the video, based on its memory. It is worth mentioning that through comprehensive evaluation, we find that existing open-source VideoLLMs suffer from severe event hallucination issues on our benchmark, which could offer promising avenues for enhancing the VideoLLM’s capabilities in the future research.
Figure 2: Illustration of examples in our EventHallusion benchmark. Questions are basically divided into two categories: Susceptibility to Language Prior and Susceptibility to Vision-Language Correlation Bias. The contents that induces VideoLLMs to generate event-related hallucinations are marked by red.
Additionally, we propose a simple yet effective method called Temporal Contrastive Decoding (TCD) to reduce the event-related hallucinations of existing VideoLLMs in a training-free manner. The core motivation of our TCD approach is to first elicit the spurious priors and then eliminate their adverse effects before making the final decisions. Technically, for each given video, we first construct its counterpart by ambiguating the temporal cues while preserving the spatial information. This counterpart disentangles the event-unrelated elements that could potentially trigger spurious prior assumptions from the original video. Therefore, we feed both the original video and its counterpart into a VideoLLM and adaptively adjust the prediction distribution of the former using the latter, leading to more reliable final judgments. As a plug-and-play approach, our TCD can consistently boost existing VideoLLM’s performances on the proposed EventHallusion as shown in Tab.2.
In summary, our contributions are threefold: (1) We propose EventHallusion, a novel video QA benchmark focused on assessing event-related hallucinations in VideoLLMs; (2) We introduce Temporal Contrastive Decoding (TCD), a simple yet effective approach that can reduce hallucinations across a wide range of VideoLLMs in a plug-and-play manner; (3) We also conduct comprehensive ablation studies to provide more in-depth analysis.
Related Works
Video Large Language Models. Large Language Models (LLMs) have achieved remarkable success in traditional NLP tasks and exhibit strong generalization to user-defined instructions, enabling them to act as versatile assistants. Building on this foundation, Multimodal Language Models (MLLMs) extend LLMs with visual inputs, facilitating unified text-vision understanding. To further incorporate video comprehension and temporality, video-supported MLLMs have emerged. Methods such as Video-ChatGPT (maaz2023videochatgpt) apply spatial-temporal pooling, while Video-LLaMA (zhang2023videollama) introduces a Video Q-Former for frame aggregation. Vista-LLaMA (ma2023vista) enhances temporal modeling via a temporal Q-Former projector, and LLaMA-VID (llamavid) proposes a dual-token paradigm to represent both context and content per frame. Although these approaches perform well on benchmarks, they still struggle with hallucinations in temporal understanding.
Video Evaluation Benchmarks. Despite the rapid progress of MLLMs, a comprehensive evaluation of their capabilities remains essential. For image MLLMs, various benchmarks have been developed to assess perception, reasoning, and cognition (fu2023mme; liu2023mmbench; li2024eyes). With the emergence of VideoLLMs, benchmarks like TempCompass (liu2024tempcompass) and FunQA (xie2023funqa) have been proposed to evaluate temporal perception and complex video comprehension.
Despite previous works sharing similarities with ours, several critical differences persist. First, our work specifically investigates event-related hallucinations in VideoLLMs, attributing their occurrence to language priors and vision-language bias inherited during pretraining. Building on this insight, we curate a benchmark with event-prone videos to facilitate targeted evaluation. Furthermore, we introduce Temporal Contrastive Decoding (TCD), a simple yet effective method to alleviate event hallucinations, demonstrating general improvements across multiple VideoLLMs.
Hallucination in MLLMs. Hallucination is a phenomenon widely observed in MLLMs, arising from factors such as imbalanced training data, weak attention to visual content, and other possible reasons. To mitigate hallucinations in image MLLMs, several strategies have been proposed. Woodpecker (yin2023woodpecker) introduces a post-hoc correction mechanism leveraging additional visual cues, while Volcano (lee2023volcano) adopts a critique-revise-decide framework for iterative self-revision. From a data perspective, LRV-Instruction (lrv_instruction) and HalluciDoctor (yu2024hallucidoctor) aim to alleviate hallucinations by constructing more balanced, high-quality instruction tuning datasets.
Another line of work addresses hallucinations during inference, where contrastive decoding (CD) has shown effectiveness without requiring additional training. VCD (leng2024vcd) mitigates statistical bias and language priors by contrasting original and distorted images during decoding. However, existing CD methods mainly target ImageLLMs, with limited exploration in the video domain. Motivated by this gap, we propose Temporal Contrastive Decoding (TCD) to reduce event hallucinations in VideoLLMs.
Figure 3: Video category and questions distribution. “Entire” implies that the entire video depicts the rare event. “Mix” implies that common and rare events interleavedly occur in the video. “Misleading” implies the video with the common event, but with its question customized to include misleading event demonstration.
EventHallusion Benchmark
In this section, we first elaborate on the data curation process of two basic categories of our EventHallusion benchmark. Afterward, we demonstrate the distribution statistics of the whole EventHallusion benchmark. Finally, we introduce the specific evaluation protocol of different question types.
Figure 4: The distribution of video length within our EventHallusion benchmark.
Data Curation
Susceptibility to language priors. In this category, we first collect videos describing the events that occur in daily life from the internet. To avoid information leakage, we blur the text captions printed on videos, thereby encouraging models to analyze events solely based on visual content. However, when annotating the questions, we explicitly include another frequently occurring event that is inconsistent with the video content to trigger the language priors within the LLM. With these “misleading triggers”, we formulate questions as a binary choice problem as shown below:
“Did the [subject] [event] happen in the video?”
Here, the combination of [subject] and [event] correspond to the “misleading triggers” aforementioned. In Fig.2, we refer to this type of sample as “Common Events with Misleading Questions”.
Figure 5: Framework of Temporal Contrastive Decoding. We construct the temporal modeling distorted video as the comparison group with fewer frames extracted and then contrast their logits to mitigate the event hallucination.
Susceptibility to vision-language correlation biases. In this category, the primary challenge lies in sourcing the appropriate video data. Since existing publicly available video datasets (e.g., ActivityNet (caba2015activitynet), (das2013thousand), (soomro2012ucf101)) collect videos from daily life, the events they include cater to the vision-language correlation biases stored in VideoLLMs, resulting in high recognition precision. Therefore, instead of refactoring existing video datasets, we meticulously collect videos from prevalent Internet video platforms and filter them with rigorous cross-validation. Specifically, we begin by manually collecting videos that depict infrequent events based on human consensus. To minimize personal bias, we engage three annotators to assess whether the events in these selected videos are indeed uncommon. Only videos that receive unanimous positive judgments from all annotators are retained as testing samples. Similar to the first category, we also employ the blurring operation on videos to prevent information leakage. We further divide these videos into two groups: (1) Videos where infrequent events occur throughout the entire duration, referred to as “Entire Rare Events” in Fig.2. (2) Videos containing a mix of both frequent and infrequent events, referred to as “Mix Common-Rare Events” in Fig.2.
For constructing questions, we design both binary and open-ended types for comprehensive evaluation. Specifically, the binary questions ask the model whether a plausible but incorrect common event occurs in the given video. The open-ended questions, on the other hand, simply prompt the model to freely describe the video content. Formally, binary and open-ended questions for the “Entire” category can be formulated as presented below:
Binary: “Did/Is the [subject] [event] in the video?”,
Open-Ended: “Please describe this video in detail.”,
For the “Mix” category, since both common and rare events can happen in the video, we reformulate the binary questions as below:
“Did the [subject] [event] in the entire video?”,
“Did any accident or anything unexpected happen in the video?”,
For the ground-truth answers to all types of questions, we meticulously verify them to ensure their correctness.
Dataset distribution statistics. In total, our EventHallusion benchmark comprises 400 videos and 711 questions. To ensure comprehensive coverage across diverse domains, these videos span seven scenarios, including Pet&Animal, Sports Competition, Food&Drink, Gym Exercises, Vehicle, Life Record, Nature, as shown in Fig.3. The distribution of video lengths in our benchmark is presented in Fig.4, showing that most videos are relatively short. This is because longer videos often involve significant scene fluctuations, which introduce additional challenges for VideoLLMs beyond the uni-modal and multi-modal biases we aim to examine in this paper.
Evaluation Protocol
Binary question answering. Following previous benchmarks (wang2024videohallucer; pope), we compare the first word of the model’s answer with the ground-truth answer to determine whether the model answered the question correctly. This approach is justified by the VideoLLM’s exceptional instruction-following capability, which typically enables it to answer binary choice questions with “yes” or “no” at the beginning. We calculate the accuracy as the evaluation metric for binary questions.
Open-ended question answering. For automatic evaluation, we feed both generated and ground-truth answers into GPT-4o(hurst2024gpt4o), using it as a knowledgeable evaluator to determine whether the model answers the question correctly following prior benchmarks (liu2024tempcompass). We also use accuracy as the evaluation metric. Detailed templates of prompting GPT-4o are provided in the supplementary materials due to page limitation.
Models Entire Mix Misleading Overall
Binary Desc. Binary Desc. Binary Binary Desc.
Open-source Models
VideoChatGPT (maaz2023videochatgpt) 14.91 5.50 56.99 3.63 21.57 36.43 4.30
VideoChat2 (li2024mvbench) 16.67 4.59 47.67 1.55 22.55 32.76 2.64
PLLaVA (xu2024pllava) 45.61 16.51 58.55 3.11 81.37 60.64 6.05
LLaMA-VID (llamavid) 30.70 16.51 73.58 7.77 43.14 54.03 10.90
LLaVA-NeXT-Video (li2024llava) 49.12 9.17 70.47 4.66 71.57 64.79 6.29
VILA (lin2024vila) 53.51 20.18 58.55 17.62 83.33 63.33 18.54
ShareGPT4Video (chen2024sharegpt4video) 11.40 0.00 67.88 5.18 6.86 49.14 9.82
Video-LLaVA (videollava) 30.70 8.26 57.51 7.25 41.18 45.97 7.62
Closed-source Models
GPT-4V (achiam2023gpt) 57.89 11.01 79.27 16.58 92.16 76.53 14.57
Gemini-1.5-Pro (team2023gemini) 61.40 49.54 83.42 39.90 96.08 80.44 43.38
GPT-4o (hurst2024gpt4o) 83.33 58.72 92.75 41.97 100.00 91.93 48.01
Table 1: Performance comparison of VideoLLMs on EventHallusion. We exhibit the accuracy of both open-source and closed-source models on the binary classification task and the detailed description matching task on our three types of videos, respectively. We highlight the best-performing results for both closed-source and open-source VideoLLMs in bold. In addition, we show the overall success rate of each model on our binary classification task that answers ”yes” or ”no” in the first word.
Temporal Contrastive Decoding
To mitigate the aforementioned hallucinations in VideoLLMs, the most straightforward solution is to enhance finetuning with additional debiased data. However, this approach is highly resource-intensive, as it requires extensive efforts to collect such videos and annotate corresponding QA pairs. Therefore, we take inspiration from (leng2024vcd) and propose our Temporal Contrastive Decoding (TCD) method, which alleviates event-oriented hallucinations in VideoLLMs by comparing the logits of the original video and its counterpart during the decoding stage.
For clarity, we first review the standard auto-regressive decoding process of a VideoLLM. Given the text instruction $x$ and the video $V$, the predicted logit $z_{t}^{ori}$ at the $t$-th timestep can be formulated as:
$z_{t}^{ori}=\operatorname{logit}_{\theta}\left(y_{t}\mid V,x,y_{<t}\right)$ (1)
where $y_{t}$ is the token to be generated at time step $t$, and $y_{<t}$ denotes the generated token sequence before time step $t$. $y_{t}$ is based on $x$, $V$, and all previously generated tokens $y_{<t}$ by selecting the token with the highest probability.
Figure 6: Performance comparison of prevalent VideoLLMs and human evaluations on object-related and event-related content recognition accuracy (i.e., our official EventHallusion evaluation) in our EventHallusion benchmark.
When constructing a contrastive counterpart, we chronologically downsample the original video as shown in the Fig.5, aiming to weaken the temporal cues while emphasizing object cues. Therefore, the contrastive counterpart’s probability prediction $z_{t}^{con}$ at timestep $t$ can be formulated as:
$z_{t}^{con}=\operatorname{logit}_{\theta}\left(y_{t}\mid S(V),x,y_{<t}\right)$ (2)
where $S(\cdot)$ represents the temporal downsampling operation. With the generated $z_{t}^{ori}$ and $z_{t}^{con}$, the modulated logit at the current timestep $t$ can be calculated as:
$z_{t}=(1+\alpha)z_{t}^{ori}-\alpha z_{t}^{con}$ (3)
where $\alpha$ is the hyperparameter to control the contrastive strength during the decoding. Since $z_{t}^{ori}$ and $z_{t}^{con}$ are not always exhibited contrastively, we eliminate the adverse effects of suppressing true positives by directly filtering implausible answers with a dynamic confidence threshold $t$:
$t=\beta*\mathrm{max}(z^{ori}_{t})$ (4)
The ultimate modulated probability $p_{t}$ can be formulated as:
[…]
$p_{t}$ $=\mathrm{softmax}(\tilde{z}_{t}).$
where $z_{t}[k]$ denotes the $k$-th dimension of $z_{t}$.
Model Decoding Entire Mix Mis. Overall
LNV (li2024llava) Default 49.12 70.47 71.57 64.79
TCD 50.88 73.58 74.51 67.48
VC2 (li2024mvbench) Default 16.67 47.67 22.55 32.76
TCD 17.54 49.74 25.49 34.72
VILA (lin2024vila) Default 53.51 58.55 83.33 63.33
TCD 50.88 66.32 84.31 66.50
Table 2: Results of TCD on EventHallusion. Default denotes the standard decoding strategy, while TCD represents our Temporal Contrastive Decoding method. We report the accuracy of the binary classification task. The best results of each category are bolded. “LNV” and “VC2” are short for “LLaVA-NeXT-Video” and “VideoChat2”, respectively. “Mis” denotes the “Misleading” category.
Experiments
In this section, we first evaluate multiple prevalent VideoLLMs on our EventHallusion, including both open-source and closed-source MLLMs. Subsequently, we examine the effectiveness of our proposed TCD and also discuss its more specific design choices.
Evaluation on EventHallusion
Performances of Human Evaluation
Firstly, to evaluate the quality of our proposed EventHallusion benchmark, we ask three humans to answer the questions and average their resulting scores as the human evaluation results. Meanwhile, to validate the value of evaluating event-oriented hallucinations for VideoLLMs, we also extracted object contents from videos in our EventHallusion benchmark and tested them using both human evaluations and two prevalent VideoLLMs. The above results are shown in Fig.6, where we have the following conclusions: (1) Although prevalent VideoLLMs perform poorly in EventHallusion (i.e., “Event” part in Fig.6), humans can correctly answer almost all questions. This demonstrates that the questions in our EventHallusion benchmark are clear enough to allow for accurate judgments. (2) When recognizing objects in videos, prevalent VideoLLMs can achieve human-level accuracies. However, they significantly lag behind humans when comprehending event-related content. This demonstrates that event-oriented hallucinations confuse VideoLLMs much more severely than object-related ones.
Performances of Prevalent VideoLLMs
Evaluation Setup. We test 8 open-source Video and 3 closed-source VideoLLMs on our EventHallusion. Open-source VideoLLMs comprise VideoChatGPT (maaz2023videochatgpt), VideoChat2 (li2024mvbench), PLLaVA (xu2024pllava), LLaMA-VID (llamavid), LLaVA-NeXT-Video (li2024llava), VILA (lin2024vila), ShareGPT4Video (chen2024sharegpt4video), and Video-LLaVA (videollava). For closed-source VideoLLMs, we select the three most popular video-supported MLLMs, namely GPT-4V (achiam2023gpt), GPT-4o(hurst2024gpt4o), and Gemini-1.5-Pro (team2023gemini), to conduct the systematic comparison of the robustness between open-source and closed-source MLLMs.
For the inference setting of each VideoLLM, we follow their original setting, including conversation mode and other hyperparameters. For the number of frames, we sample at least 8 frames for each VideoLLM to ensure the complete preservation of event information. In addition, to ensure the stability and reproduction of our results, we adopt the greedy search as the default decoding strategy for all open-source models. For the closed-source models, we sample 8 frames.
Method Entire Mix Misleading Overall
Baseline 49.12 70.47 71.57 64.79
TCD 50.88 73.58 74.50 67.48
TCD+noise 51.75 73.58 74.50 67.73
Table 3: Ablation study of different strategies when constructing the contrastive counterpart of the original video. The best results of each category are bolded.
Performance Analysis. We exhibit the performance of each VideoLLM in Tab 1 on the binary classification and detailed description matching task. As the results indicate, open-source VideoLLMs exhibit a consistent weakness in event hallucinations. In contrast, closed-source VideoLLMs demonstrate greater robustness against event hallucinations across all categories. A deeper analysis is provided below: (1) In evaluating the susceptibility of VideoLLMs to vision-language correlation biases (i.e., “Entire” and “Mix” in Tab.1), most VideoLLMs perform worse when confronted with videos containing a mix of common and rare events, compared to those containing only rare events. This result indicates that frequent events can distract the model’s attention, imposing higher challenges to the model. (2) In evaluating the susceptibility of VideoLLMs to language priors (i.e., “Misleading” in Tab.1), closed-source models demonstrate overwhelming advantages compared to open-source models, suggesting that these closed-source VideoLLMs are built upon language foundation models that are significantly more powerful than their open-source counterparts.
TCD Evaluation
VideoLLM Baselines.
To evaluate the effectiveness of our TCD in addressing event hallucinations, we adopt LLaVA-NeXT-Video (li2024llava), VideoChat2 (li2024mvbench), and VILA (lin2024vila) as the foundation backbones. LLaVA-NeXT-Video extends the ability of LLaVA (liu2024visual) based on the projection layer design, while VideoChat2 adopts the Q-Former (li2023blip) to bridge the gap between videos and LLM. VILA, on the other hand, employs enhanced visual-language pre-training for better video comprehension. Based on these MLLMs, we explore the performance of TCD across different VideoLLM architectures and designs on EventHallusion.
Figure 7: Effects of $\alpha$ and $\beta$ in our TCD.
Implementation Details.
We adopt different numbers of extracted frames to construct the original and distorted visual input: 32 and 8 frames for LLaVA-NeXT-Video, 12 and 8 frames for VILA, and 16 and 4 frames for VideoChat2. We use greedy search as the basic decoding paradigm.
Effectiveness of TCD
We apply our proposed TCD method on three prevalent VideoLLMs, namely LLaVA-NeXT (li2024llava), VideoChat2 (li2024mvbench), and VILA (lin2024vila), and evaluate their performances on EventHallusion. As shown in Tab.2, equipping with TCD, three VideoLLMs achieve a consistent performance boost. This demonstrates that TCD can be seamlessly integrated with versatile VideoLLMs and reduce their event-related hallucinations without bells and whistles.
Explorations on Contrastive Sample Constructions
We further explore more alternative strategies to construct the contrastive counterpart in our TCD method. Recalling that our TCD aims at ambiguating the temporal cues of the original video, we further explore the effects of blurring the spatial cues by adding gaussian noises to video frames when constructing the contrastive counterpart. As shown in Tab.3, introducing Gaussian noises on top of TCD can further boost the overall performance. However, the extent of introducing Gaussian noise should be deliberately adjusted to prevent degradation in practice. Therefore, we do not choose to apply this strategy in our TCD method.
Frames Entire Mix Misleading Overall
1 48.26 68.39 74.51 64.30
4 50.00 73.06 76.47 67.48
8 50.88 73.58 74.51 67.48
16 50.88 73.06 74.51 66.75
Table 4: Ablation study of extracted video frames when constructing the contrastive video.
Ablations on Hyper-parameters of TCD
We also ablate the effects of $\alpha$ and $\beta$ in our TCD method. As shown in Fig.7, we have the following observations: (1) With $\alpha$ increases from 0.25 to 0.5, TCD enjoys the benefits of the contrastive video’s modulation. However, further increasing $\alpha$ to 1.0 causes TCD to degrade significantly, suggesting that excessively amplifying the influence of original samples will cause the model to decode primarily according to the original distribution, rendering TCD ineffective. (2) Increasing $\beta$ from 0.2 to 0.5 enhances the model’s performance, while further decreasing or increasing $\beta$ does not result in performance fluctuations. This demonstrates that our TCD is relatively robust to the variation in $\beta$ values.
Effect of Sampled Frames in Distorted Video.
We further explore the effects of the temporal downsampling ratio when constructing the contrastive video in our TCD. Specifically, we adopt LLaVA-NeXT-Video(li2024llava) as the backbone model with 32 frames sampled for the original video input and downsample the number of sampled frames to construct the contrastive video. This experiment is conducted on the binary classification task of EventHallusion. From Tab.4, we have the following observations: (1) Extreme temporal downsampling (i.e., 1 frame) can lead to severe performance degradation of our TCD approach. This phenomenon may arise because single-frame input significantly diminishes spatial cues, leading to distorted prediction logits during contrastive learning. (2) A low temporal downsampling rate (e.g., 16 frames) performs worse than higher rates (e.g., 4 and 8 frames). This indicates that narrowing the gap between the contrastive video and the original one weakens the advantages of our TCD approach.
Conclusion
In this paper, we present EventHallusion, a novel benchmark designed to systematically evaluate the event-related hallucination issues in existing VideoLLMs. To address these hallucinations, we propose a simple yet effective Temporal Contrastive Decoding (TCD) method, which can be seamlessly integrated into prevalent VideoLLMs in a plug-and-play fashion. Extensive experimental results demonstrate the challenging nature of EventHallusion and the effectiveness of the proposed TCD approach.
Appendix A Comparison with Other Video Hallucination Benchmarks.
We exhibit the comparison between HallusionBench, VideoHallucer, and our EventHallusion in Figure 8. As shown in Figure 8, compared to HallusionBench and VideoHallucer, which emphasize assessing the models’ sensitivity to events’ chronological orders, our proposed EventHallusion focuses on analyzing the event-related language prior and vision-language correlation bias encapsulated in existing VideoLLMs.
Appendix B Automatic Evaluation for Detailed Description Matching
We apply GPT-4o to evaluate the detailed description matching task of our Entire, Mix, and Misleading categories, respectively. For Entire, we prompt GPT-4o to judge if VideoLLMs’ detailed description of the video is consistent with the ground-truth rare event. And for Mix, we prompt GPT4o to focus on the occurrence of the rare-common mixed event. Finally, for Misleading, we prompt GPT4o to assess the consistency of the ground-truth event and search for any unrelated events. In addition, we set up the default system prompt to better guide GPT-4o’s evaluation, which is shown in Figure 9. The detailed prompts are shown in Figure  10, 11, 12, respectively.
Figure 8: Comparison between HallusionBench, VideoHallucer, and our EventHallusion. We use green to highlight the content of the original video/question, and red to denote refactored video/question content to assess the model’s event-related hallucinations.
Figure 9: System prompt used in all automatic evaluation.
Appendix C Qualitative Examples
Data examples of EventHallusion. We present examples of Entire, Mix, and Misleading in Figure 13, 14, 15, incorporating responses of VideoLLMs. As shown in these examples, VideoLLMs are prone to generate hallucinated answers under all of our settings, indicating their deficiency in addressing event hallucinations.
Examples of GPT-4o’s Evaluation. We present examples of GPT-4o’s evaluation of detailed description matching task on Entire, Mix, and Misleading respectively in Figure 16, 17, 18. As shown in these examples, GPT-4o has the capacity to judge the accuracy of video events.
Examples of TCD. We present examples of our TCD in mitigating video event hallucinations. Specifically, we exhibit the generated response with and without TCD for LLaVA-NeXT-Video and VideoChat2 on our EventHallusion in Figure 19, 20, and 21.
Figure 10: Prompt used for detailed description matching in Entire.
Figure 11: Prompt used for detailed description matching in Mix.
Figure 12: Prompt used for detailed description matching in Misleading.
Figure 13: Examples of Entire with responses of VideoLLMs.
Figure 14: Examples of Mix with responses of VideoLLMs.
Figure 15: Examples of Misleading with responses of VideoLLMs.
Figure 16: Examples of the evaluation of GPT-4o on the detailed description of VideoLLMs of Entire.
Figure 17: Examples of the evaluation of GPT-4o on the detailed description of VideoLLMs of Mix.
Figure 18: Examples of the evaluation of GPT-4o on the detailed description of VideoLLMs of Misleading.
Figure 19: Examples of TCD with LLaVA-NeXT-Video on our EventHallusion.
Figure 20: Examples of TCD with LLaVA-NeXT-Video on our EventHallusion.
Figure 21: Examples of TCD with VideoChat2 on our EventHallusion.